# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v1

@@ Abstract
To support extremely high data rates, reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is considered to be a promising technology for future sixth-generation networks. However, due to the typical employment of hybrid beamforming architecture in THz systems, as well as the passive nature of RIS which lacks the capability to process pilot signals, obtaining channel state information (CSI) is facing significant challenges. To accurately estimate the cascaded channel, we propose a novel low-complexity channel estimation scheme, which includes three steps. Specifically, we first estimate full CSI within a small subset of subcarriers (SCs). Then, we acquire angular information at base station and RIS based on the full CSI. Finally, we derive spatial directions and recover full-CSI for the remaining SCs. Theoretical analysis and simulation results demonstrate that the proposed scheme can achieve superior performance in terms of normalized mean-square-error and exhibit a lower computational complexity compared with the existing algorithms.
@@ I Introduction
Envisioned as a cornerstone technology to meet the requirements of future 6G network, terahertz (THz) communication holds immense potential for a myriad of applications within the field of Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1]. However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [2]. One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [3]. To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide LoS paths by intelligently manipulating the propagation environment [4]. RIS can dynamically adjust reflection properties of a large number of passive reflecting elements to optimize signal transmission and reception [5]. By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems. However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), of which acquisition can be challenging. Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
@@ I-A Prior Work
Over the past few years, numerous research works have been dedicated to explore channel estimation in RIS-assisted high frequency band systems [6, 7, 8, 9, 10, 11]. Specifically, a sparse representation of RIS-assisted millimeter (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [6]. In [7], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel. To reduce computational complexity, [8] proposed a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation. The authors in [9] proposed a two-phase channel estimation strategy to reduce pilot overhead. [10] investigated cooperative-beam-training based channel estimation procedure. In [11], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
@@
The above solutions have effectively addressed the channel estimation problem of mmWave and narrowband THz systems with utilization of the angular domain sparsity. However, THz communication typically adopts a hybrid beamforming architecture to reduce hardware costs[12]. This architecture introduces the beam split effect, where path components split into distinct spatial directions at different SC frequencies due to the wider bandwidth of THz signals and the enormous number of RIS elements [13]. This phenomenon leads to a considerable degradation in achievable data rates, counteracting the benefits of RIS-assisted THz systems.
@@
To address this issue, beam split effect for RIS-assisted THz systems have been investigated in [14, 15, 16]. In particular, [14] introduced a beam-split-aware-orthogonal-matching-pursuit (BSA-OMP) approach to automatically mitigate beam split. In [15], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator. The authors in [16] proposed a robust regularized sensing-beam-split-OMP based scheme to estimate beam split affected RIS-assisted cascaded channel. The above works have achieved accurate channel estimation under beam split effect with the identical core idea of mitigating beam split effect by a well-designed dictionary covering the entire beamspace. However, these algorithms will need to be executed at each SC to ensure estimation accuracy, which may introduce additional unaffordable computations. To further reduce the complexity, literatures [17, 18, 19] have investigated methods by prioritizing the solution of physical directions. Specifically, the authors in [17] regarded channel estimation problem as a block-sparse recovery problem. [18] proposed a beam split pattern detection based wideband channel estimation scheme where the physcical DoAs are prioritized for determination. In [19], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals(AoAs) and angle-of-departure(AoDs) are determined and refined to reconstructed the wideband channel with beam split effect. However, in RIS-assisted THz systems, due to the additional angular information and the coupling of different physical directions introduced by RIS, the aforementioned solutions might be challenging to maintain their low-complexity advantages. Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
@@ I-B Main Contributions
To address the issue discussed above, we presented some preliminary results to estimate the RIS-assisted cascaded channel [20]. In this paper, we significantly expand the previous work and propose a low-complexity channel estimation scheme for RIS-assisted wideband THz systems with beam split. Our main contributions are summarized as follows:
  • We focus on the channel estimation in a general RIS-assisted wideband THz communication scenario with consideration of the beam split effect. In this scenario, we propose a novel low-complexity channel estimation scheme with three steps. To recover the full CSI within a small subset of SCs, we propose a cyclic-beam-split generalized-approximate-message-passing (CBS-GAMP) approach in Step $1$. To reduce computational complexity, we propose a novel beam split angle decoupling (BSAD) scheme where we propose energy-maximum (EnM) and double-sensing-multiple-signal-classification (DS-MUSIC) algorithm to acquire angular information in Step $2$ and we utilize least-squares (LS) solution to obtain the rest angle-excluded information in Step $3$.
  • For the CBS-GAMP approach, we design simplified CBS dictionaries to provide comprehensive coverage of the entire beam space and construct the sparse representation of cascaded channel. We employ this algorithm with expectation-maximum (EM) extensions to acquire full CSI within a small subset of SCs, without the need of prior knowledge of channel distribution and signal-to-noise-ratio (SNR).
  • For the BSAD scheme, we extend the full CSI acquired by CBS-GAMP approach within a small subset of SCs to all the SCs. Specifically, we propose a novel EnM algorithm to estimate the directions-of-departure (DoDs) at BS. Additionally, we design an extended-log multiple-signal-classification pseudo-spectrum (EL-MUSIC-PS) and propose a novel DS-MUSIC algorithm to estimate the coupled directions at RIS. Subsequently, we incorporate the spatial directions derived from the estimated angular information to the unknown CSI and reconstruct the RIS-assisted cascaded channel by a simple LS solution.
The remainder of this paper is organized as follows. In Section II, we explain the RIS-assisted wideband THz system model with beam split. In Section III, we propose a novel CBS-GAMP based full CSI estimation approach. In Section IV, we propose a further simplified BSAD scheme. In Section V, we discuss the numerical simulations and then draw conclusions in Section VI.
@@
Notations: Boldface lower-case and capital letters represent column vectors and matrices, respectively. $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively. $\mathbb{E}$ and $\mathrm{Var}$ represent statistical expectation and variance, respectively. $\mathbb{N}$, $\mathbb{C}$ and $\mathbb{R}$ denote the set of integers, complex numbers and real numbers, respectively. $\mathbf{I}_{N}$ is $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace. $\otimes$ denote Kronecker product, *** and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
II RIS-assisted wideband THz system model
In this section, we first present system model and downlink channel estimation protocol for the considered RIS-assisted wideband THz system with beam split effect. Then, we formulate the RIS-assisted THz channel model and further elaborate beam split effect.
II-A RIS-Assisted Wideband THz System Model
@@
Figure 1: RIS-assisted THz communication system with beam split.
Figure 2: Downlink channel estimation protocol and frame structure.
@@
We consider a RIS-assisted wideband THz MIMO system with hybrid analog/digital beamforming architecture over $M$ subcarriers (SCs), as depicted in Fig. 1. Assuming that the BS with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}(N_{\mathrm{RF}}<N_{\mathrm{T}})$ ratio-chains are deployed to serve $K$ single-antenna UEs. The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS. Besides, the OFDM system is considered with the central frequency $f_{c}$ and bandwidth $B$. The overall downlink channel estimation protocol is shown in Fig. 2. We assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [9]. For each coherence block, channel estimation is firstly performed during $T$ subframes, and the downlink data transmission is then carried out during the remaining $D-T$ subframes. Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots.
Assuming that the direct BS-UE links are obstructed by obstacles [17], the downlink channel $\mathbf{H}_{k}^{\mathrm{RIS}}\left[m\right]\in\mathbb{C}^{1\times N_{\mathrm{T}}}$ from BS to UE-$k$ at the $m$-th SC can be expressed as
$\begin{split}\mathbf{H}_{k}^{\mathrm{RIS}}\left[m\right]&=\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}\right)\mathbf{G}\left[m\right]\\ &=\boldsymbol{\theta}\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right],\end{split}$ (1)
where $\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right]\triangleq\mathbf{G}\left[m\right]\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ is the RIS-assisted cascaded channel and $\boldsymbol{\theta}=\left[e^{j\theta_{1}},\cdots,e^{j\theta_{N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS.
@@
As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[12]. As mentioned above, we consider that $P$ frequency-dependent baseband beamformers $\mathbf{W}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}\triangleq\mathrm{Diag}\left\{s_{1},\cdots,s_{P}\right\}\in\mathbb{C}^{P\times P}$ per time slot. Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain augment in each user direction, wherein each entry is subject to the unit-modulus constraint, i.e., $|\left[\mathbf{W}^{\mathrm{RF}}\right]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ as $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$. Define $\mathbf{W}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}[m]=\left[\mathbf{w}_{1}[m],...,\mathbf{w}_{P}[m]\right]$. At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by UE-$k$ at the $m$-th SC is denoted as
@@
$\begin{split}y_{k,t,p}[m]&=\mathbf{\theta}_{t}\mathbf{H}_{k}^{\mathrm{cas}}[m]\mathbf{w}_{t,p}[m]s_{t,p}[m]+n_{k,t,p}[m],\end{split}$ (2)
@@
where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN). Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t,p}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}\times P}}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
@@
$\tilde{\mathbf{y}}_{k,t}[m]=\boldsymbol{\theta}_{t}\mathbf{H}_{k}^{\mathrm{cas}}[m]\mathbf{R}_{t}[m]+\tilde{\mathbf{n}}_{k,t}[m].$ (3)
@@
We adopt the widely used Saleh-Valenzuela channel model [21] for BS-RIS channel $\mathbf{G}[m]$ and RIS-UE channel $\mathbf{h}_{k}[m]$. The BS-RIS channel $\mathbf{G}[m]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ at the $m$-th SC is expressed as

### v2

# Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split

@@ Abstract
Reconfigurable intelligent surface (RIS)-assisted terahertz (THz) communication is emerging as a key technology to support ultra-high data rates in future sixth-generation networks. However, the acquisition of accurate channel state information (CSI) in such systems is challenging due to the passive nature of RIS and the hybrid beamforming architecture typically employed in THz systems. To address these challenges, we propose a novel low-complexity two-phase channel estimation scheme for RIS-assisted THz systems with beam split effect. In the proposed scheme, we first estimate the full CSI over a small subset of subcarriers, then extract angular information at both the base station and RIS. Subsequently, we recover the full CSI across remaining subcarriers by determining the corresponding spatial directions and angle-excluded coefficients. Theoretical analysis and simulation results demonstrate that the proposed method achieves superior performance in terms of normalized mean-square error while significantly reducing computational complexity compared to existing algorithms.
@@ I Introduction
Envisioned as a cornerstone technology to meet the requirements of future 6G networks, terahertz (THz) communication holds immense potential for a myriad of applications within the field of the Internet of Things, ranging from intelligent industrial manufacturing to digital twinning technologies and immersive experiences like holographic telepresence and the metaverse [1, 2, 3]. However, THz communication systems encounter various challenges, including high propagation loss, limited coverage, and susceptibility to atmospheric absorption and scattering [4, 5]. One notable drawback is the severe signal attenuation experienced by transmitted signals, which often necessitate a line-of-sight (LoS) link to maintain communication quality [6]. To address these challenges and unlock the full potential of THz communication, reconfigurable intelligent surface (RIS) has emerged to provide LoS paths by intelligently manipulating the propagation environment [7]. RIS can dynamically adjust reflection properties of a large number of passive reflecting elements to optimize signal transmission and reception [8]. By strategically controlling the phase shifts of these reflecting elements, RIS can effectively enhance the signal strength, mitigate interference, and improve the overall performance of THz communication systems. However, due to the passive nature of RIS, which lacks signal processing capabilities, the performance enhancement enabled by RIS relies significantly on accurate channel state information (CSI), the acquisition of which can be challenging. Therefore, investigating an efficient channel estimation scheme for RIS-assisted THz communication systems is of great importance.
@@ I-A Prior Work
Over the past few years, numerous research works have been dedicated to exploring channel estimation in RIS-assisted high frequency band systems [9, 10, 11, 12, 13, 14]. Specifically, a sparse representation of RIS-assisted millimeter wave (mmWave) channel has been established, on which compressed sensing based techniques can be implemented for CSI acquisition [9]. In [10], a low-overhead channel estimation method has been proposed based on the double-structure characteristic of RIS-assisted angular cascaded channel. To reduce computational complexity, [11] proposed a non-iterative two-stage framework based on direction-of-arrival (DoA) estimation. The authors in [12] proposed a two-phase channel estimation strategy to reduce pilot overhead. [13] investigated cooperative-beam-training based channel estimation procedure. In [14], a sparse Bayesian learning based channel estimation approach has been developed to compensate the severe off-grid issue.
@@
The above solutions have effectively addressed the channel estimation problem of narrowband systems with utilization of the angular domain sparsity. However, THz communication typically adopts a hybrid beamforming architecture to reduce hardware costs[15]. When considering this architecture, the beam split effect becomes inevitable. This effect occurs due to the wider bandwidth of THz signals and the enormous number of RIS elements, which causes path components to split into distinct spatial directions at different subcarrier frequencies [16]. This phenomenon leads to a considerable degradation in achievable data rates, which counteracts the benefits of RIS-assisted THz systems.
@@
To address this issue, beam split effect in RIS-assisted THz systems has been investigated in [17, 18, 19]. In particular, [17] introduced a beam-split-aware orthogonal matching pursuit (BSA-OMP) approach to automatically mitigate beam split. In [18], a low-complexity beam squint mitigation scheme is proposed based on generalized-OMP estimator. The authors in [19] proposed a robust regularized sensing-beam-split-OMP-based scheme to estimate beam split affected RIS-assisted cascaded channel. These works primarily focus on mitigating beam split effects within the context of cascaded channels using well-designed dictionaries covering the entire beamspace. However, an alternative approach is to focus on the physical directions of the channel components to reduce computational complexity [20, 21, 22]. Specifically, the authors in [20] regarded channel estimation problem as a block-sparse recovery problem. [21] proposed a beam split pattern detection based wideband channel estimation scheme where the physical DoAs are prioritized for determination. In [22], a three-stage wideband channel estimation scheme is proposed where angle-of-arrivals (AoAs) and angle-of-departure (AoDs) are determined and refined to reconstruct the wideband channel with beam split effect. Although these algorithms achieve accurate channel estimation, they require execution at each SC or simultaneous processing of all SCs, which may introduce additional unaffordable computations. Moreover, the passive nature of RIS complicates the precise acquisition of angles at the RIS, which often results in the availability of only coupled angles. Therefore, this motivates the development of a low-complexity channel estimation scheme in RIS-assisted wideband THz systems with beam split effect.
@@ I-B Main Contributions
To address the issue discussed above, we presented some preliminary results on estimating the RIS-assisted cascaded channel [23]. In this paper, we significantly expand on the previous work and propose a two-phase efficient channel estimation scheme for RIS-assisted wideband THz systems with beam split. Our main contributions are summarized as follows:
  • We propose a novel low-complexity channel estimation protocol for general RIS-assisted wideband THz communication scenarios with beam split effect, as depicted in Fig. 1. The channel estimation scheme is divided into two phases: full CSI estimation for two SCs in Phase I and remaining SCs estimation in Phase II. By employing information from only two SCs, this protocol can significantly reduce the computational complexity while maintaining complete and accurate channel estimation.
  • In Phase I, we propose a two-step method to acquire full CSI for the selected two SCs. Specifically, in Step 1, we propose a simplified cyclic beam split (CBS) dictionary to reduce the size from the scale of $Q^{2}$ to a linear scale of $Q$. Using this dictionary, we reformulate the problem as a generalized sparse recovery problem. Moreover, we discuss the selection of different sparse recovery algorithms. In Step 2, we first design a heuristic DoD estimation method at BS according to the estimated cascaded channel obtained in Step 1. We then reformulate the coupled angles at RIS as a typical DoA estimation problem and design a double sensing-multiple signal classification (DS-MUSIC) algorithm by integrating the angular information from the two selected SCs. We also provide insights into the selection of two SCs. In Phase II, given angular information from Phase I, we employ a simple least squares (LS) algorithm to estimate the channels for the remaining SCs.
  • Through the analysis of algorithm complexity, our proposed protocol significantly reduces the computational complexity compared to the existing sparse recovery solutions. Additionally, numerical results demonstrate that the NMSE performance of our proposed channel estimation scheme closely approaches that of oracle-LS method. By focusing on a two-step estimation process for two SCs and applying a simplified LS algorithm for the remaining SCs, the proposed method can achieve both reduction of computational complexity and acquisition of accurate CSI.
Figure 1: Two-phase channel estimation protocol and frame structure.
The remainder of this paper is organized as follows. In Section II, we first describe the RIS-assisted wideband THz communication setup and the beam split effect. In Section III, we describe our proposed channel estimation protocol. Section V provides an analysis of the computational complexity of our proposed scheme, with a comparison to existing methods. Finally, we conclude the paper in Section VI.
@@
Notations: Boldface lower-case and capital letters represent column vectors and matrices, respectively. $\left(\cdot\right)^{*}$, $\left(\cdot\right)^{\mathsf{T}}$, $\left(\cdot\right)^{\mathsf{H}}$ and $\left(\cdot\right)^{\dagger}$ denote conjugate, transpose, transpose-conjugate and pseudo-inverse operation, respectively. $\mathbb{E}$ and $\mathrm{Var}$ represent statistical expectation and variance, respectively. $\mathbb{N}$, $\mathbb{C}$ and $\mathbb{R}$ denote the set of integers, complex numbers and real numbers, respectively. $\mathbf{I}_{N}$ is $N\times N$ size identity matrix and $\mathrm{tr}$ is the trace. $\otimes$ denote Kronecker product, * and $\bullet$ denote Khatri-Rao product and transposed Khatri-Rao product, respectively.
II System and Channel Model
II-A Signal and Channel Model
@@
Figure 2: RIS-assisted THz communication system with beam split effect.
@@
We consider a RIS-assisted wideband THz MIMO system employing hybrid beamforming architecture that integrates both analog and baseband processing, as depicted in Fig. 2. Assuming that the BS equipped with $N_{\mathrm{T}}$ antennas and $N_{\mathrm{RF}}$ (where $N_{\mathrm{RF}}<N_{\mathrm{T}}$) RF chains is deployed to serve $K$ single-antenna UEs. The RIS employs a uniform linear array (ULA) comprised of $N_{\mathrm{R}}$ reflecting elements, which is manipulated by a smart controller operated by the BS. Besides, the OFDM system is considered with $M$ subcarriers, central frequency $f_{c}$ and bandwidth $B$.
In this paper, we assume that the channel is quasi-static, i.e., the channel remains approximately constant in a channel coherence block with $D$ subframes [12]. For each coherence block, channel estimation is firstly performed during $T$ subframes, and downlink data transmission is then carried out during the remaining $D-T$ subframes. Moreover, each subframe of the downlink channel estimation period is divided into $P$ time slots, with $P$ baseband beamformers operating alternatively across different time slots. Additionally, we assume that the direct BS-UE links are blocked by obstacles [20]. Therefore, we mainly focus on the downlink channel estimation of the cascaded BS-RIS-UE channel.
@@
As a cost-performance trade-off, hybrid beamforming architecture is typically employed in THz MIMO systems[15]. In our considered system, we assume that $P$ frequency-dependent baseband beamformers $\mathbf{W}_{t}^{\mathrm{BB}}[m]=\left[\mathbf{w}^{\mathrm{BB}}_{t,1}[m],\ldots,\mathbf{w}^{\mathrm{BB}}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{RF}}\times P}$ are employed to process $P$ orthogonal pilot signals $\mathbf{S}_{t}\triangleq\mathrm{Dia{\tiny}g}\left\{s_{t,1},\cdots,s_{t,P}\right\}\in\mathbb{C}^{P\times P}$ per time slot. Meanwhile, a frequency-independent analog beamformer $\mathbf{W}^{\mathrm{RF}}\in\mathbb{C}^{N_{\mathrm{T}}\times N_{\mathrm{RF}}}$ is employed to achieve gain improvement in the direction of RIS with each entry is subject to the unit-modulus constraint, i.e., $|[\mathbf{W}^{\mathrm{RF}}]_{i,j}|=1/\sqrt{N_{\mathrm{T}}}$ for $i=1,\ldots,N_{\mathrm{T}}$ and $j=1,\ldots,N_{\mathrm{RF}}$. We assume that the analog beamformer remains unchanged over multiple channel coherence blocks due to the relatively static nature of the RIS physical location compared to the dynamic channel coefficients. Therefore, we define $\mathbf{W}_{t}[m]\triangleq\mathbf{W}^{\mathrm{RF}}\mathbf{W}_{t}^{\mathrm{BB}}[m]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$ as the hybrid beamformer which consists of $P$ beamformer vectors, i.e., $\mathbf{W}_{t}[m]=\left[\mathbf{w}_{t,1}[m],...,\mathbf{w}_{t,P}[m]\right]$. Let $\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}$ and $\mathbf{h}_{k}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}}$ denote the channel from BS to RIS and RIS to $k$-th UE at the $m$-th SC, respectively. At time slot $p$ of subframe $t$, the received measurement $y_{k,t,p}[m]$ by $k$-th UE at the $m$-th SC can be expressed as
@@
$\begin{split}y_{k,t,p}[m]\!&=\!\mathbf{h}_{k}\left[m\right]\mathrm{Diag}\left(\boldsymbol{\theta}_{t}\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m]\\ &=\!\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{w}_{t,p}[m]s_{t,p}[m]\!+\!n_{k,t,p}[m],\end{split}$ (1)
where $n_{k,t,p}[m]\sim\mathcal{CN}\left(0,2\varepsilon\right)$ is the complex additive white Gaussian noise (AWGN). Besides, $\boldsymbol{\theta}_{t}=\left[e^{j\theta_{t,1}},\cdots,e^{j\theta_{t,N_{\mathrm{R}}}}\right]$ is the phase-shift matrix of RIS at subframe $t$ with a constant reflection amplitude constraint [24].
@@
Let $\mathbf{r}_{t,p}[m]\triangleq\mathbf{w}_{t,p}[m]s_{t,p}[m]\in\mathbb{C}^{N_{\mathrm{T}}}$ denote the processed transmitted signal and combining the processed signals during $P$ successive time slots $\mathbf{R}_{t}[m]\triangleq\left[\mathbf{r}_{t,1}[m],...,\mathbf{r}_{t,P}[m]\right]\in\mathbb{C}^{N_{\mathrm{T}}\times P}$, the combined measurement vector $\tilde{\mathbf{y}}_{k,t}[m]\in\mathbb{C}^{1\times P}$ is expressed as
@@
$\tilde{\mathbf{y}}_{k,t}[m]=\boldsymbol{\theta}_{t}\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\!\mathbf{G}\left[m\right]\mathbf{R}_{t}[m]+\tilde{\mathbf{n}}_{k,t}[m],$ (2)
where $\tilde{\mathbf{n}}_{k,t}[m]\triangleq\left[n_{k,t,1}[m],...,n_{k,t,P}[m]\right]\in\mathbb{C}^{1\times P}$ denotes the combined AWGN. Therefore, the cascaded channels can be expressed as [25]
$\mathbf{H}_{k}^{\mathrm{cas}}\left[m\right]=\mathrm{Diag}\left(\mathbf{h}_{k}\left[m\right]\right)\mathbf{G}\left[m\right]\in\mathbb{C}^{N_{\mathrm{R}}\times N_{\mathrm{T}}}.$ (3)
@@
We adopt the widely used Saleh-Valenzuela channel model for both BS-RIS channel and RIS-UE channel [26]. The BS-RIS channel $\mathbf{G}[m]$ at the $m$-th SC is expressed as